# reptheorem[1]

### v1

@@ 1 Introduction
The problem of inverse reinforcement learning (IRL) can be informally characterized as follows. Given observations of an expert acting optimally with respect to an unknown reward in an MDP, we aim to recover a reward for which the expert is optimal. While early inquiries can be traced back to optimal control theory Kalman1964 and to econometrics Rust1992 , IRL was first introduced to the machine learning community by Russell Russell1998 . The motivation behind IRL is two-fold: First, IRL is a powerful tool for imitation learning, where the goal is to recover the expert’s policy from a dataset of expert demonstrations. This is particularly useful in scenarios such as autonomous driving pomerleau1988alvinn where it is easy to collect expert demonstrations. Second, compared to other imitation learning methods that only recover a policy, IRL comes with the advantage that recovering a reward provides a more interpretable and transferable description of the task, as the reward can be potentially used to learn optimal policies in a new environment.
@@ 4.1 Reward convergence
Moreover, to recover a reward for which the expert is $(\varepsilon+\varepsilon_{\text{real}})$-optimal we require the length of the expert trajectories to be $H=\mathcal{O}(\log\varepsilon^{-1})$ and we need an expected number of $\mathcal{O}(1/\varepsilon^{2})$ samples from the MDP.
@@
$\mathbb{E}\left[\max_{\pi}L(\pi,w^{t})-L(\pi^{t},w^{t})\right]\leq\mathcal{O}\left({1/\sqrt{T}}\right),$ (13)
where the expectation is taken with respect to all the randomness in Algorithm 1. The above lemma shows that if we control the changes of the reward $r^{t}$ by setting the reward learning rate to $\eta_{w}=\Theta(1/\sqrt{T})$, the value of the policy $\pi^{t}$ converges to the optimal value under $r^{t}$ at the same speed $\mathcal{O}(1/\sqrt{T})$. We provide the proof in Appendix A.
@@ 4.2 Policy convergence
We define $\pi$ to be $\varepsilon$-close to the expert policy if the total variation metric described above is bounded by $\varepsilon$. To ensure convergence in the policy, we need the following additional assumption. {asm}[] The state occupancy measure $\nu^{\pi^{E}}(s)$ generated by the expert satisfies $\vartheta^{E}:=\min_{s\in\mathcal{S}}\nu^{\pi^{E}}(s)>0$. The above assumption ensures that the expert sufficiently explores the state space of the MDP. Similar to Assumption 4.1, it is satisfied when the initial distribution $\nu_{0}$ is bounded away from zero. Under Assumption 16, we have the following convergence guarantee for the optimal policy corresponding to the recovered reward $\bar{r}$.
@@
Suppose Assumptions 4.1, 4.2, and 16 hold, and let $\eta_{w}=\frac{1-\gamma}{\sqrt{kT}\|\phi\|_{\infty}}$. Algorithm 1 requires an expected number of $\mathcal{O}(1/\varepsilon^{4})$ samples to ensure that the optimal policy corresponding to the recovered reward $\bar{r}$ is $(\varepsilon+\sqrt{\varepsilon_{\text{real}}})$-close to the expert policy.
@@
where the equality $(i)$ follows from the soft suboptimality Lemma (Mei2020, , Lemma 26), $(ii)$ holds by Pinsker’s inequality, and $(iii)$ by Assumption 16. Therefore, to ensure $\max_{s}\|\pi^{E}(\cdot|s)-\pi^{*}_{r^{\bar{w}}}(\cdot|s)\|_{\text{TV}}\leq\varepsilon+\sqrt{\varepsilon_{\text{real}}}$, we need $J^{*}_{\bar{r}}-J^{\pi^{E}}_{\bar{r}}$ to be upper bounded by $\Omega\left(\varepsilon^{2}+\varepsilon_{\text{real}}\right)$. By Theorem 4.3, we need an expected number of $\mathcal{O}(1/\varepsilon^{4})$ samples in total. ∎
@@
As shown by Proposition 4.5, convergence in the total variation metric (16) is stronger than convergence in the integral probability metric (22). This is because the total variation metric directly measures the difference between policies, while the integral probability metric measures the difference in their unregularized expected value, ignoring the entropy regularization values. In the above example, the expert policy $\pi^{E}$ has maximum entropy and is therefore realizable with $\varepsilon_{\text{real}}=0$. Therefore, Algorithm 1 is guaranteed to recover a reward with a corresponding optimal policy that is $\varepsilon$-close to the expert policy in the total variation metric. However, in the case of a large realizability error $\varepsilon_{\text{real}}$, Corollary 4.4 fails to provide such strong guarantees. Hence, compared with Syed2007 ; viano2022proximal ; Shani2021 ; Liu2022 , we can get stronger convergence guarantees for the optimal corresponding to the recovered reward if the expert is realizable with a small realizability error $\varepsilon_{\text{real}}$.
@@ Appendix A Proof of main lemmas
For estimator $\hat{\sigma}^{E}$, we have, for all $w\in W$,
@@
For reward gradient estimator ^$\hat{\nabla}_{w}L(\pi,w)$, we have
@@
From [1][Assumption 6.3], we have $\mathbb{E}\left[\hat{Q}^{\pi}_{r}(s,a)\right]={Q}^{\pi}_{r}(s,a)$ and $(s_{H},a_{H})$ with $H\sim\operatorname{Geom}(1-\gamma)$ is sampled from the state-action occupancy measure $\mu^{\pi}$. Therefore, we have
@@
$\leq$ $\frac{2\eta_{w}\|\phi\|_{\infty}}{(1-\gamma)^{2}}-\left(\min_{s}\frac{\nu^{\pi^{t+1}}(s)}{\nu^{\pi_{r^{t}}^{*}}(s)}\right)\frac{\tau}{1-\gamma}\mathbb{E}\left[\mathbb{E}_{s\sim\nu^{\pi_{r^{t}}^{*}}}\text{KL}(\pi^{t}(\cdot|s)||\pi^{t+1}(\cdot|s))\right]$ (39)

### v3

@@ 1 Introduction
The problem of inverse reinforcement learning (IRL) can be informally characterized as follows. Given observations of an expert acting optimally with respect to an unknown reward in a Markov decision process (MDP), we aim to recover a reward for which the expert is optimal. While early inquiries can be traced back to optimal control theory Kalman1964 and to econometrics Rust1992 , IRL was first introduced to the machine learning community by Russell Russell1998 . The motivation behind IRL is two-fold: First, IRL is a powerful tool for imitation learning, where the goal is to recover the expert’s policy from a dataset of expert demonstrations. This is particularly useful in scenarios such as autonomous driving pomerleau1988alvinn where it is easy to collect expert demonstrations. Second, compared to other imitation learning methods that only recover a policy, IRL comes with the advantage that recovering a reward provides a more interpretable and transferable description of the task, as the reward can be potentially used to learn optimal policies in a new environment.
@@ 4.1 Reward convergence
Here, the expectation is taken with respect to all the randomness in Algorithm 1. Moreover, to recover a reward for which the expert is $(\varepsilon+\varepsilon_{\text{real}})$-optimal we require the length of the expert trajectories to be $H=\mathcal{O}(\log\varepsilon^{-1})$ and we need an expected number of $\mathcal{O}(1/\varepsilon^{2})$ samples from the MDP.
@@
$\mathbb{E}\left[\max_{\pi}L(\pi,w^{t})-L(\pi^{t},w^{t})\right]\leq\mathcal{O}\left({1/\sqrt{T}}\right).$ (13)
The above lemma shows that if we control the changes of the reward $r^{t}$ by setting the reward learning rate to $\eta_{w}=\Theta(1/\sqrt{T})$, the value of the policy $\pi^{t}$ converges to the optimal value under $r^{t}$ at the same speed $\mathcal{O}(1/\sqrt{T})$. We provide the proof in Appendix A.
@@ 4.2 Policy convergence
We define $\pi$ to be $\varepsilon$-close to the expert policy if the total variation metric described above is bounded by $\varepsilon$. To ensure convergence in the policy, we need the following additional assumption.
Assumption 4.4.
The state occupancy measure $\nu^{\pi^{E}}(s)$ generated by the expert satisfies $\vartheta^{E}:=\min_{s\in\mathcal{S}}\nu^{\pi^{E}}(s)>0$.
The above assumption ensures that the expert sufficiently explores the state space of the MDP. Similar to Assumption 4.1, it is satisfied when the initial distribution $\nu_{0}$ is bounded away from zero. Under Assumption 4.4, we have the following convergence guarantee for the optimal policy corresponding to the recovered reward $\bar{r}$.
@@
Suppose Assumptions 4.1, 4.2, and 4.4 hold, and let $\eta_{w}=\frac{1-\gamma}{\sqrt{kT}\|\phi\|_{\infty}}$. Algorithm 1 requires an expected number of $\mathcal{O}(1/\varepsilon^{4})$ samples to ensure that the optimal policy corresponding to the recovered reward $\bar{r}$ is $(\varepsilon+\sqrt{\varepsilon_{\text{real}}})$-close to the expert policy.
@@
where the equality $(i)$ follows from the soft suboptimality Lemma (Mei2020, , Lemma 26), $(ii)$ holds by Pinsker’s inequality, and $(iii)$ by Assumption 4.4. Therefore, to ensure $\max_{s}\|\pi^{E}(\cdot|s)-\pi^{*}_{r^{\bar{w}}}(\cdot|s)\|_{\text{TV}}\leq\varepsilon+\sqrt{\varepsilon_{\text{real}}}$, we need $J^{*}_{\bar{r}}-J^{\pi^{E}}_{\bar{r}}$ to be upper bounded by $\Omega\left(\varepsilon^{2}+\varepsilon_{\text{real}}\right)$. By Theorem 4.3, we need an expected number of $\mathcal{O}(1/\varepsilon^{4})$ samples in total. ∎
@@
As shown by Proposition 4.6, convergence in the total variation metric (16) is stronger than convergence in the integral probability metric (22). This is because the total variation metric directly measures the difference between policies, while the integral probability metric measures the difference in their unregularized expected value, ignoring the entropy regularization values. In the above example, the expert policy $\pi^{E}$ has maximum entropy and is therefore realizable with $\varepsilon_{\text{real}}=0$. Therefore, Algorithm 1 is guaranteed to recover a reward with a corresponding optimal policy that is $\varepsilon$-close to the expert policy in the total variation metric. However, in the case of a large realizability error $\varepsilon_{\text{real}}$, Corollary 4.5 fails to provide such strong guarantees. Hence, compared with Syed2007 ; viano2022proximal ; Shani2021 ; Liu2022 , we can get stronger convergence guarantees for the optimal policy corresponding to the recovered reward if the expert is realizable with a small realizability error $\varepsilon_{\text{real}}$.
@@ Appendix A Proof of main lemmas
For the estimator $\hat{\sigma}^{E}$, we have, for all $w\in W$,
@@
For the reward gradient estimator ^$\hat{\nabla}_{w}L(\pi,w)$, we have
@@
Here, the expectations are with respect to the randomness of the corresponding sampling strategies.
@@
From [1, Assumption 6.3], we have $\mathbb{E}\left[\hat{Q}^{\pi}_{r}(s,a)\right]={Q}^{\pi}_{r}(s,a)$ and $(s_{H},a_{H})$ with $H\sim\operatorname{Geom}(1-\gamma)$ is sampled from the state-action occupancy measure $\mu^{\pi}$. Therefore, we have
@@
$\leq$ $\frac{2\eta_{w}\|\phi\|_{\infty}}{(1-\gamma)^{2}}-\frac{\tau}{1-\gamma}\mathbb{E}\left[\left({\min_{s}\frac{\nu^{\pi^{t+1}}(s)}{\nu^{\pi_{r^{t}}^{*}}(s)}}\right)\mathbb{E}_{s\sim\nu^{\pi_{r^{t}}^{*}}}\text{KL}(\pi^{t}(\cdot|s)||\pi^{t+1}(\cdot|s))\right]$ (39)